OpenStudy (sharkfinsoup):

State the degree: rst+r^2s^2+r^3s^2 would the answer be the highest exponent (3) or all exponents added together (9)?

OpenStudy (sharkfinsoup):

I went on mathwarehouse.com and they said that you just choose the largest one.

OpenStudy (sharkfinsoup):

Ok so 3 was wrong and this was their reasoning "The correct answer is: 5. Remember that the degree of a polynomial is equal to the highest degree of any of its terms." so you have to treat rs as a pair?

Nnesha (nnesha):

well not for this polynomial bec there are 3 variables so in this case the highest degree would be the sum of all the exponents in each term \[r^1 s^1 t^1 = 1+1+1=3\] so the first term has a degree of 3

OpenStudy (sharkfinsoup):

oh ok the next one would be 4 and the last term would be 5 which is the highest one

Nnesha (nnesha):

correct

OpenStudy (sharkfinsoup):

ok thank you

Nnesha (nnesha):

yw :=))

OpenStudy (jerry45):

no